You purchase 14 call option contracts with a strike price of $80 and a premium of $1.80. Assume the stock price at expiration is $92.00.

a. What is your dollar profit?

b. What is your dollar profit if the stock price is $77.95? (A negative value should be indicated by a minus sign. Do not round intermediate calculations.)

f the stock price is $77.95, the call is , so the dollar profit is   
Answer #1

(a) Strike Price = $ 80 and Price at Expiry = $ 92, It is assumed that each call option is for 100 stocks

Number of Options Bought = 14 and Payoff Per Option = (92 - 80) x 100 = $ 1200

Total Pay Off = 12 x 14 x 100 = $ 16800 and Premium per Option = $ 1.8

Total Premium Paid = 1.8 x 14 x100 = $ 2520

Net Payoff = $ Profit = (16800 - 2520) = $ 14280

(b) If the stock price is $ 77.95, the call option will not be exercised as the strike price is above the underlying asset's price and hence option payoff = $ 0

Net Payoff = $ Profit = 0 - 25 20 = - $ 2520 (negative sign indicates $ loss)